OpenStudy (anonymous):

wo numbers add to 272 and the first is 254 bigger than the second. What are the two numbers?

OpenStudy (aum):

Let the two numbers be x and y. x + y = 272 x - y = 254 Solve for x and y.

OpenStudy (anonymous):

how do you solve for x and y with a problem like this

OpenStudy (anonymous):

@CollegeCountryGuy @aum Saying it that way may be a bit confusing. You need to substitute a term. We know one term is 254 larger than the second. So we know: \[x+(x+254)=272\] Now solve for x.

OpenStudy (aum):

Add the two equations and that will get rid of y and you can solve for x. The put x back in one of the equations and solve for y.

OpenStudy (anonymous):

im still confused

OpenStudy (aum):

x + y = 272 x - y = 254 add: 2x = 526 divide both sides by 2: x = 263 Put x = 263 in the second equation: 263 - y = 254 y = 263 - 254 = 9 The two numbers are 263 and 9.

OpenStudy (anonymous):

Two numbers add to 473 and the first is 77 bigger than the second. What are the two numbers?

OpenStudy (anonymous):

@CollegeCountryGuy You should really be able to solve that now. It is the same procedure. If you have math anxiety or something similar you might try doing your work on light purple paper. Studies have shown it to be helpful for people with dyslexia and dyscalculia.